# ColPali: Efficient Document Retrieval with Vision Language Models

### v3

@@ B.1 Codebase
B.2 Pairwise CE loss
Our in-batch contrastive loss $\mathcal{L}$ is defined as the softmaxed cross-entropy of the positive scores $s_{k}^{+}=\text{LI}\left(d_{k},q_{k}\right)$ w.r.t. to the maximal negative scores $s_{k}^{-}=\max\limits_{l,l\neq k}\hskip 8.53581pt\text{LI}\left(q_{k},p_{l}\right)$.
For numerical stability, we reformulate the loss with the softplus function, leading to:
$\mathcal{L}=\frac{1}{b}\sum_{k=1}^{b}\texttt{softplus}\left(s_{k}^{-}-s_{k}^{+}\right)$ (2)
B.3 Hyperparameters
Hyperparameters are tuned on a validation split composed of $2\%$ of the training dataset. We find bi-encoder methods to be more sensible to learning rate variations than late interaction-based models and achieve the best performance for all models with a learning rate of $55e-$ 5. We experiment with LoRA rank and $\alpha$ values and do not notice particular improvements past $r=\alpha=32$. Per-device batch sizes are kept small due to long sequence lengths that complicate scaling past $b=4$. Simulating larger batch sizes for in-batch negative sampling should enable even better results. We find the best results with global batch size $b=32$ for 1 epoch on our training set.
B.4 Embedding size
Minimizing storage footprint can be essential to industrial retrieval systems if databases contain millions of documents. With this criterion in view, we have compared the embedding sizes of the models in our study. As shown in Table 3, ColPali’s embedding size is an order of magnitude larger than BM25 and two orders of magnitude larger than BGE-M3. However, this study is limited to the naive method of storing ColPali’s multi-vector embeddings. In practical scenarios, using cluster centroids can reduce the size of ColPali multi-vector embeddings by up to an order of magnitude (Santhanam et al., 2022) and make it a competitive retrieval system.
@@ B.4 Embedding size
ColPali (float16) 256
Table 3: Comparison of the embedding sizes for the DocVQA test set from ViDoRe w.r.t. different retrieval models. The lower the size the smaller the storage footprint of the model. The mean ± std size is given for the sparse embeddings.
B.5 Latency computations
All latency computations are done on a NVIDIA L4 GPU. Queries are encoded independently (batch size of 1) to simulate online querying, and pages are encoded with a batch size of 4 for PaliGemma derived models, and 8 for BGE-M3. Reported times include image and text processing time before the model forward pass, as well as query-to-index matching times. We note an interesting feature of ColPali is that all documents have the same sequence length, leading to prior knowledge of runtime and memory consumptions. Query latency experiments are averaged over 1000 queries, and indexing times are measured for a 100 page document. Per page time is obtained by diving total time by 100, corresponding to inverse page throughput.
B.6 Captioning
@@
Figure 5: Example from the "Energy" test set.
@@
Figure 6: Example from the "Government Reports" test set.
Caption: The image shows a table titled "System of Record" which outlines the different types of documents or records maintained across various systems or departments within an organization related to project management and construction. The rows list documents like project plans, budgets, schedules, contracts, purchase orders, invoices, change requests, bid submissions, drawings, manuals, meeting minutes, and reports. The columns indicate the system or department responsible for maintaining each record, such as County Servers, Project View, OnBase, CGI Advantage Financial System, and Purchasing Department. The table uses "W" and "T" markers to denote which system or department serves as the primary source (writer) or storage location (trailer) for each type of document.
Appendix C More similarity maps
In Figure 7, ColPali assigns a high similarity to all patches with the word "Kazakhstan" when given the token <_Kazakhstan>. Moreover, our model seems to exhibit world knowledge capabilities as the patch around the word "Kashagan" - an offshore oil field in Kazakhstan - also shows a high similarity score. On the other hand, in Figure 8, we observe that ColPali is also capable of complex image understanding. Not only are the patches containing the word "formulations" highly similar to the query token _formula, but so is the upper-left molecule structure.
Figure 8: Similarity of the image patches w.r.t. the underlined token in the user query. This example is from the Healthcare Industry test set.
It is also interesting to highlight that both similarity maps showcase a few white patches with high similarity scores. This behavior might first seem surprising as the white patches should not carry a meaningful signal from the original images. We believe the vectors associated with these patches share a similar role with the ViT registers (Darcet et al., 2023), i.e. these patches were repurposed for internal computations and stored the global information from the whole image.
Appendix D Additional results
D.1 Other Metrics
@@
Unstructured Text only
@@ D.2 Model Variants
ColPali (No Mem. Tokens) 80.4 53.2 82.4 77.4 65.7 63.4 97.0 89.9 93.6 92.4 79.6
ColPali (Best) 79.1 54.4 81.8 83.9 65.8 73.2 96.2 91.0 92.7 94.4 81.3
Table 5: Evaluation of some "negative results" and ablations on ViDoRe; ColPali for reference. Results are presented using NDCG@5 metrics. Text-only metrics are not computed for benchmarks with only visual elements.
Appendix E ViDoRe examples

### v4

@@ B.1 Codebase
B.2 Hyperparameters
Hyperparameters are tuned on a validation split composed of $2\%$ of the training dataset. We find bi-encoder methods to be more sensible to learning rate variations than late interaction-based models and achieve the best performance for all models with a learning rate of $55e-$ 5. We experiment with LoRA rank and $\alpha$ values and do not notice particular improvements past $r=\alpha=32$. Per-device batch sizes are kept small due to long sequence lengths that complicate scaling past $b=4$. We simulate larger batch sizes with multi-GPU training and train with a total batch size $b=32$ with no accumulation, for 1 epoch on our training set.
B.3 Embedding size
Minimizing storage footprint can be essential to industrial retrieval systems if databases contain millions of documents. With this criterion in view, we have compared the embedding sizes of the models in our study. As shown in Table 4, ColPali’s embedding size is an order of magnitude larger than BM25 and two orders of magnitude larger than BGE-M3. However, in practical scenarios, pooling multi-vector embeddings by centroid cluster, or quantizing embeddings to binary representations  [^20] https://blog.vespa.ai/scaling-colpali-to-billions/ can reduce storage costs by two orders of magnitude (Santhanam et al., 2022) with minimal performance hits, and make storage costs competitive with other systems.
@@ B.3 Embedding size
ColPali (float16) 257.5
Table 4: Comparison of the embedding sizes for the DocVQA test set from ViDoRe w.r.t. different retrieval models. The mean ± std size is given for the sparse embeddings. In general multiple vectors (2-5) per page are used for BGE-M3 and BM25.
B.4 Latency computations
To ensure comparison fairness, the latencies of the different retrieval systems shown in Figure 2 are measured on the same g2-standard-8 GCP VM with a NVIDIA L4 GPU. Document pages are embedded using the highest settings of Unstructured with captioning (see subsection 3.2). SigLIP and ColPali are both loaded with bfloat16 parameter dtypes. The reported times in Table 5 are the average per-page latencies for each indexing operation on 1000 randomly chosen documents across all splits of the ViDoRe benchmark test set. A batch size of $8$ was used for the BGE-M3 model used with Unstructured, and a batch size of $4$ was used for SigLIP and ColPali.
Indexing operation Latency (s)
Unstructured SigLIP ColPali
Layout detection 0.81 NA NA
OCR 2.67 NA NA
Captioning 3.71 NA NA
Page encoding 0.03 0.12 0.39
Total 7.22 0.12 0.39
Table 5: Page-level latencies for document indexing using various retrieval systems. SigLIP and ColPali are much faster than Unstructured because they don’t require the layout detection, OCR, and captioning operations.
B.5 Captioning
@@
Figure 4: Example from the “Energy” test set.
@@
Figure 5: Example from the “Government Reports” test set.
Caption: The image shows a table titled “System of Record” which outlines the different types of documents or records maintained across various systems or departments within an organization related to project management and construction. The rows list documents like project plans, budgets, schedules, contracts, purchase orders, invoices, change requests, bid submissions, drawings, manuals, meeting minutes, and reports. The columns indicate the system or department responsible for maintaining each record, such as County Servers, Project View, OnBase, CGI Advantage Financial System, and Purchasing Department. The table uses ”W” and ”T” markers to denote which system or department serves as the primary source (writer) or storage location (trailer) for each type of document.
Appendix C Additional results
C.1 Other Metrics
@@
Unstructured text-only
@@ C.2 Model Variants
ColPali (No Q.A. Tokens) 80.4 53.2 82.4 77.4 65.7 63.4 97.0 89.9 93.6 92.4 79.6
ColPali (Docmatix) 71.3 48.0 80.0 83.9 59.1 73.8 95.7 93.8 92.5 93.1 79.1
ColPali (224) 71.0 37.4 62.3 65.7 28.6 20.4 65.7 66.8 73.9 73.0 56.5
ColPali (Vision Trained) 78.8 53.9 81.3 81.7 64.4 70.6 95.3 91.7 93.5 94.7 80.6
ColPali (No Pairwise) 79.0 53.0 82.1 85.3 63.2 66.2 94.9 88.9 92.7 92.1 79.7
ColPali (+TabFQuAD training) 77.6 54.7 82.6 86.5 65.4 73.9 94.8 92.4 94.2 94.8 81.7
ColIdefics2 (64) 73.6 48.0 82.4 81.6 63.0 57.2 95.5 86.9 86.6 91.2 76.6
ColQwen2 (768) 86.4 56.2 89.8 88.7 75.2 85.7 98.8 94.8 93.6 97.3 86.6
ColPali (Reference: 448) 79.1 54.4 81.8 83.9 65.8 73.2 96.2 91.0 92.7 94.4 81.3
Table 7: Benchmark scores for the “negative results” and various ablations on ViDoRe; ColPali for reference. Results are presented using nDCG@5 metrics. Text-only metrics are not computed for benchmarks with only visual elements.
Appendix D More similarity maps
In Figure 6, ColPali assigns a high similarity to all patches with the word “Kazakhstan” when given the token <_Kazakhstan>. Moreover, our model seems to exhibit world knowledge capabilities as the patch around the word ”Kashagan”—an offshore oil field in Kazakhstan—also shows a high similarity score.
It is also interesting to highlight that both this similarity map and the one displayed in Figure 3 (right) showcase a few white patches with high similarity scores. This behavior might first seem surprising as the white patches should not carry a meaningful signal from the original images. We believe the vectors associated with these patches share a similar role with the ViT registers (Darcet et al., 2023), i.e. these patches were repurposed for internal computations and stored the global information from the whole image.
Appendix E Model glossary
SigLIP
SigLIP (Sigmoid Loss for Language Image Pre-Training) builds upon CLIP (Contrastive Language-Image Pretraining)—a foundational model that aligns images and text by maximizing the similarity between correct image-text pairs while minimizing it for incorrect ones, leveraging a contrastive loss (Zhai et al., 2023). Unlike CLIP (Radford et al., 2021), which applies the softmax function to the logits, SigLIP uses the sigmoid activation function. This innovation eliminates the need for a global view of all pairwise similarities between images and texts within a batch, enabling more flexible batch size scaling (up to 1M items per batch, with an effective optimal batch size of 32k). This approach allows SigLIP to achieve state-of-the-art performance in zero-shot image classification tasks.
PaliGemma
PaliGemma is a 3B-parameter vision-language model. It integrates the SigLIP vision encoder with a Gemma-2B language decoder, connected via a multimodal linear projection layer (Lucas Beyer* et al., 2024). The model processes images by segmenting them into a fixed number of Vision Transformer (Dosovitskiy et al., 2020) tokens, which are prepended to an optional text prompt.
A distinguishing feature of PaliGemma is its operation as a Prefix-Language Model (Prefix-LM). This design ensures full attention between image tokens and the user-provided input (prefix) while generating outputs auto-regressively (suffix). This architecture allows image tokens to access the task-specific query during processing, facilitating more effective task-dependent reasoning.
PaliGemma was trained in four stages: unimodal pretraining with existing components, extended multimodal pretraining, short high-resolution pretraining, and task-specific fine-tuning.
ColBERT
ColBERT (Contextualized Late Interaction over BERT) is a retrieval model designed to balance speed and effectiveness in information retrieval tasks (Khattab & Zaharia, 2020). Traditional retrieval models are typically categorized based on their type of interaction: either processing queries and documents independently for efficiency (bi-encoders) or jointly to capture rich contextual relationships (cross-encoders). ColBERT combines the advantages of both approaches through a novel late interaction mechanism.
Queries and documents are encoded separately using BERT, enabling offline pre-computation of document representations for scalability. Instead of pooling embeddings into a single vector, ColBERT retains token-level embeddings and employs a MaxSim operator to compute fine-grained similarity scores. For each query token, the model determines the maximum similarity with document tokens, summing these scores to compute relevance.
This architecture preserves the contextual richness of deep language models while significantly improving computational efficiency. By delaying the interaction step, ColBERT supports vector similarity indexing, facilitating end-to-end retrieval from large collections without prohibitive costs. Empirical evaluations on passage search datasets demonstrate that ColBERT achieves competitive effectiveness compared to existing BERT-based models (Devlin et al., 2018), while executing queries orders of magnitude faster and with drastically reduced computational requirements.
Appendix F Examples from the ViDoRe benchmark